OpenStudy (owen3):

Epsilon Delta proof (shudders). Show that n^1/n -> 1 as n-> ∞

OpenStudy (owen3):

Let \(\rm \large h_n = n^{1/n} -1 \) Then\[\rm \large n = ( 1 + h_n ) ^n \\ \large = 1 + nh_n + n(n-1) \frac{(h_n)^2}{2!} ... + (h_n)^n > 1 + nh_n \] From which it follows \[ \rm \large n > 1 + nh_n \\ \large \implies h_n < \frac{n-1}{n} \\ \large \implies h_n < 1 - \frac 1 n < 1 + \frac 1 n ~~~~ \text{ since } n > 0 \] Then \[ \large \rm | h_n - 0 |= h_n < 1+ \frac 1 n < \epsilon~ \\ \large \rm \text{ if } n \ge N = \frac{1}{ \epsilon -1 }+1 \] We can assume that \( \epsilon \neq 1 \) above

OpenStudy (jango_in_dtown):

There will be a problem... since epsilon can be as small as we please

OpenStudy (owen3):

the inequality 1 - 1/n < e did not produce n > f(e)

OpenStudy (eliesaab):

1+1/n -> 1 and cannot be made close to zero

OpenStudy (jango_in_dtown):

right...

OpenStudy (owen3):

So i can't salvage this proof?

OpenStudy (jango_in_dtown):

Yeah, the method I used, thats correct.. elisaab told the correct thing, 1+1/n is always greater than 1

OpenStudy (owen3):

ok this is a bad proof since e can be less than 1 but 1 + 1/n is always greater than 1

OpenStudy (jango_in_dtown):

ANy ways, you tried to give an alternative proof.. Its always good to try something new

OpenStudy (owen3):

1 - 1/n < e 1 - e < 1/n n < 1 / (1 - e ) there might be a way to make n > something

OpenStudy (owen3):

Oh, in the proof we need \( \rm h_n\) to get as close to zero (the limit) as possible? That is not possible with \( \large \rm 1 - \frac 1 n \) or \( \large \rm 1 + \frac 1 n \)

OpenStudy (owen3):

Or in other words, we need to bound 1\( \pm \) 1\n by epsilon, but epsilon goes to zero

OpenStudy (jango_in_dtown):

right.

OpenStudy (owen3):

and since n goes to infinity, that can't go to zero

OpenStudy (owen3):

I like how in your proof you have | h(n) - 0 | < e in general it seems like you can always do that i.e. if lim f(n) = L , then h(n) = f(n) - L and then find the limit of this. This seems to simplify things because you can concentrate just on the epsilon